OpenStudy (anonymous):

solve sqrt x-3≥7 x-3 is squar rooted

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@hedyeh99

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

x less than or equal to 3

jimthompson5910 (jim_thompson5910):

no, not sure how you got that

OpenStudy (anonymous):

whats my answeR?

jimthompson5910 (jim_thompson5910):

what do you get when you solve \[\Large \sqrt{x-3} = 7\]

OpenStudy (anonymous):

x≤52

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

close, but somehow you flipped the sign

jimthompson5910 (jim_thompson5910):

it should be \[\Large x \ge 52\]